OpenStudy (anonymous):

Which of the numbers below could be terms in the sequence an = 7n + 4? A. 46 B. 32 C. 24 D. 36

OpenStudy (anonymous):

B and A??

OpenStudy (anonymous):

C. ...... 24-4=20 not a multiple of 7 so no B. ....... 32-4=28 28/7= 7 so yes A. .... 46-4=42 42/7=6 so yess D. .....36-4= 32 not a multiple of 7 @Romero

OpenStudy (anonymous):

28/7=4* sorry

myininaya (myininaya):

Yes @SantanaG you are right!

OpenStudy (anonymous):

Thank you @myininaya :)